Find the range of the function:
f(x)= x^2-12x+46

Answer by lwsshak3(11628)

Find the range of the function:
f(x)= x^2-12x+46
complete the square
y=(x^2-12x+36)+46-36
y=(x-6)^2+10
This is an equation of a parabola that opens upwards with a vertex at (6, 10), that is, it has a minimum of 10.
The range of the function, therefore, is (10,∞)